# S-box Security Analysis of NIST Lightweight Cryptography Candidates: A Critical Empirical Study

### v1

@@ 3.3 Differential Cryptanalysis
Differential cryptanalysis is a technique that relies on exploiting chosen plaintext pairs and resulting ciphertext pairs, in order to gain the information of secret key. The pairs of plaintext and ciphertext are referred as differentials. And these differentials when combined for multiple rounds become differential characteristics of a cipher under study [47]. The probabilities of occurrence of certain output (ciphertext) differences corresponding fixed input (plaintext) differences are computed in the form of DDT table. A low value of differential uniformity, high differential branch number, high value of propagation characteristics, and no undisturbed bits are a requirement for an sbox to resist against differential cryptanalysis and its variant attacks.
@@
[…]
@@
When an affine transformation is applied to $S$, DDT preserves the maximum entry and weight of coefficients in rows and columns, however their position is subject to a change [103, 104].
Definition 3.4
@@
For an sbox to resist differential cryptanalysis a smaller value of DU is required. If an affine transformation is applied to $S$. DU stays invariant [104].
@@
An sbox having DBN greater than $2$ increases the chances of having more active sboxes in iterative ciphers. DBN does not stay invariant under affine equivalence [98].
@@
$PC(l)$ of $S$ is not variant when an affine transformation is applied [87]. An sbox satisfies PC when $ACT_{max}=0$ For an sbox to satisfy SAC, complying with PC is a pre-requisite [74]. Sboxes having low order $PC(l)$ have weak diffusion characteristics [73].
@@
For a specific nonzero input difference $\triangle_{\varkappa}\in F_{2}^{n}$ of an sbox $S:F_{2}^{n}\rightarrow F_{2}^{m}$, some of the bits in the corresponding output difference $\triangle_{\rho}$ remain unchanged, these specific bits are referred to as undisturbed bits (UDB) [108].
@@
=$\triangle_{\flat}=\{\triangle_{\rho}=\rho_{j}\in F_{2}^{m}\ |Pr_{S}(\triangle_{\varkappa}\rightarrow\triangle_{\rho})>0\}$ (24)
Here $j=(0,1,2,....m-1)$, if $\rho_{j}=d$, when $d\in F_{2}$ is fixed and $\triangle_{\rho}\in\triangle_{\flat}$ then $S$ has UDBs. UDB is invariant under affine transformation. An sbox should have no UDB in order to avoid the construction of truncated differentials. Currently, there is no theoretically proven upper bound on UDB.
@@ 3.4 Boomerang Cryptanalysis
Boomerang attack is a technique of block cipher cryptanalysis based on the similar principle as differential cryptanlysis. However, instead of drawing one differential characteristic for the entire cipher, boomerang attack divides the cipher into two sub ciphers and generates two differentials characteristics also known as boomerang distinguishers. A low value of boomerang uniformity makes an sbox resistant towards boomerang attacks.
@@
The highest value in BCT is known as the boomerang uniformity (BU); where $[\triangle_{\varkappa},\nabla_{\flat}]\neq 0$ [55].
@@
BCT and BU of an sbox are invariant under affine transformation [74, 55].
@@ 3.5 Differential-Linear Cryptanalysis
Bar-On et al [57] proposed Differential Linear Connectivity Table (DLCT) in order to study the dependencies between two parts of the cipher so that better differential and linear attack characteristics could be drawn. The DLCT of $S$ is a $2^{n}\times 2^{m}$ table, where rows contain input differences and columns contain output bit masks.
For $\triangle_{\varkappa}\in F_{2}^{n}$ and $\rho\in F_{2}^{m}$, the DLCT entry $(\triangle_{\varkappa},\rho)$ can be computed mathematically as follows:-
$\mathsf{DLCT}_{S}(\triangle_{\varkappa},\rho)=\#\{\varkappa\in F_{2}^{n}|\ \rho\cdot S(\varkappa)=\rho\cdot S(\varkappa+\triangle_{\varkappa})\}-2^{n-1}$ (27)
@@
$\Gamma_{S}=\{\mathsf{DLCT}_{S}(\triangle_{\varkappa},\rho),\triangle_{\varkappa}\in F_{2}^{n}\backslash\{0\}\mbox{ and }\rho\in F_{2}^{m}\backslash\{0\}\}$ (28)
@@
Under affine equivalence, DLCT and DLU are invariant [111, 78]. In order to resist against differential linear attack, an sbox should have low DLU.
@@ 3.6 Algebraic Cryptanalysis
The basic idea behind algebraic attacks involves converting the entire block cipher round function into a system of equations and then trying to solve these equations in order to deduce the secret key. Both, non-linear and linear components are transformed into polynomials. Unlike the linear component, the number of variables for nonlinear layer increases after every next round. Optimal sboxes are better resistant towards algebraic crypltanalysis if they have a high value of algebraic degree and algebraic immunity, and high degree of interpolation polynomial.
@@
Where the coefficients $c_{i}\in F_{2}$ form the elements of the truth table for ANF of $f(\varkappa)$.
An n-bit vectorial boolean function $S:F_{2}^{n}\rightarrow F_{2}^{m}$ has $f_{j}=f_{1},f_{2},f_{3},.....f_{n}$ coordinate functions and each can be uniquely represented in the form of ANF.
@@
The AD is an affine invariant property of any mapping $S$ [101, 74]. Sboxes having a low AD are vulnerable to higher order differential attacks. AD has an inverse relation with CI, the higher is the order of CI, the lower is AD.
@@
In 1996, Jakobsen and Knudsen introduced interpolation attack on block ciphers[112]. In this attack the round function or sboxes are represented in the form of algebraic expressions. In either case, Lagrange interpolation technique is used to evaluate the algebraic expression. The Lagrange polynomial is constructed using as many plaintext/ ciphertext pairs as the number of unknown coefficients in the algebraic representation. Attack complexity depends on number of terms or degree of the resulting polynomial expression.
For any boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ with input $\varkappa_{i}=\{\varkappa_{1},.....,\varkappa_{n}\}$ and output $\flat_{i}=\{\flat_{1},.....,\flat_{n}\}$ where $i=1,......,n$, we find the Interpolation Polynomial (IP) $P_{i}(\varkappa)$ as follows:
@@
IP is invariant under affine equivalence [113], and a low degree of IP makes any sbox susceptible to interpolation attack.
@@ 3.7 Side Channel Analysis
Attacks on cryptographic implementations have evolved significantly since Kocher’s first attempt to use physical leakages from a cryptosystem to extract sensitive information. Due to the increasing threat, designers are always trying to come up with efficient ways to resist against side channel attacks [114]. Starting from the design of components such as sboxes with good cryptographic properties to resist side channels and moving up to the overall cipher level by employing countermeasures such as boolean or polynomial masking. Properties such as DPA SNR and transparency order are nowadays actively used to determine an sbox’s resilience against DPA attacks. However, striking a perfect balance between the metrics related to traditional cryptanalysis and side channel analysis is an open research area. Currently, designers try to achieve cryptographic profiles that are well suited with the requirements of a particular infrastructure or standardization process (as in NIST LWC). In ideal terms, for an sbox to be resistant towards attacks such as DPA, it needs to posses a low DPA SNR and transparency order.
@@
If an sbox has small value of TO, it is considered highly resistant against DPA attacks [61]. When affine transformation is applied to an sbox, TO does not stay invariant [81].
4 Results and Analysis
Table 5 presents the sbox analysis results of Romulus, ASCON/ ISAP, Elephant, Photon-Beetle, and GIFT-COFB. Results of ASCON and ISAP are listed in the same column, since both have same sbox.
Diffusion Metrics
@@ 4 Results and Analysis
If we look at the basic avalanche characteristics of all these sboxes, they satisfy the criteria for balancedness, ensuring equal number of $1$’s and 00’s in their coordinate and component functions. This helps reduce the chances of getting any high probability linear approximations, in turn assuring the resistance of these sboxes towards linear cryptanalysis[79]. The sboxes under analysis are bijective, so they satisfy the criteria of permutation. From the bounds discussed in Table 3, we have developed an understanding that ideally the sboxes should have a full cycle length for order of permutation. Presumably, sboxes having short iterative periods are vulnerable to attacks so designers take advantage of the affine equivalence criteria and try to improve the order of permutation of sboxes using affine transformation [50, 51, 83, 84]. Out of the six, Ascon/ ISAP sbox has a FP at $\varkappa=255$. Photon-Beetle/ Elephant/ GIFT-COFB, each has one OFPs at $\varkappa=14/8/3$. Sboxes having FPs can alter the overall diffusion power that an iterative cipher has to offer and make it vulnerable to invariant subspace attacks, it can be fixed using affine equivalence. However, in practice, many notable cipher designs avoid the presence of FPs, due to the reason of achieving good results in other affine variant properties. BIC value for all sboxes is $1.0000$ which indicates that the input and output variables have a high correlation amongst them. Furthermore, none of sboxes satisfy SAC. Both these findings indicate that an attacker can gain statistical information about the sboxes and use it to exploit the entire cryptosystem [85, 115]. AI values for all the sboxes meet UB, and the values for that of SSI seem a bit better but do not satisfy the LB requirement. Hence, results from both the properties suggest that sboxes have weak diffusion characteristics.
Linear Metrics
@@
Security towards linear cryptanalysis attacks is ensured when an sbox has a low LAP and high (close to maximum) degree of NL. All $6$ sboxes under analysis have obtained the desired bounds on both properties. LBN of all $6$ sboxes is closer to the LB, which is opposite to the ideal requirements. However, if the design requires it, LBN can be improved using affine transformations. The presence of nonzero LS makes it possible for the cipher to leak significant amount of information if specific group(s) of bits in the input are flipped. All the sboxes are first order correlation immune, which makes them secure against zero-correlation attacks.
Differential and Boomerang Metrics
@@
Almost Perfect Nonlinear (APN) sboxes, in odd dimensions are differentially $2$ uniform [51]. The only odd dimension case in sboxes under study is of Ascon/ ISAP, and since its has a DU: $8$, we can classify it as not an APN. All these sboxes have DU values within the desired bounds. DBN of Elephant and Photon-Beetle satisfies the UB, Ascon/ ISAP is closer to the UB and for Romulus and GIFT-COFB DBN value stands on the LB. A good DBN ensures more number of active sboxes across multiple rounds in block ciphers. As discussed earlier that for an sbox to satisfy SAC one of the pre-requisites is compliance with PC. In Table 5 we can see that all sboxes have PC $=0$, which again points towards the weak avalanche profile.
UDB are detected in all $6$ sboxes. In [108] authors have extensively described how the presence of UDB in an sbox impacts other cryptographic properties, such as the occurrence of LS, as we can see in Table 5. C. Tezcan used these UDB and LS to devise 4 and 5 round impossible, truncated and improbable differentials, and break 5 out of 12 Ascon rounds [116]. In regards to boomerang cryptanalysis, the BU for all sboxes is attaining the UB, which is opposite to the required bounds in order to be secure against Boomerang attacks.
Differential-Linear Metrics
@@
Combined attacks make it easier for an attacker to break a cipher by defining short differential and linear characteristics, as in the case of differential-linear cryptanalysis. Hence, attaining the LB on DLU is crucial to gain resistance against these attacks. In case of sboxes under analysis, we can observe from Table 5, that the DLU is a little way up from the LB, hence paving a way for differential-linear attack with a reduced complexity.
Algebraic Metrics
@@
Previously, we have talked about the inverse relation between CI and AD, and we observed that all $6$ sboxes have a CI $=0$. So given the inverse relation, AD of all sboxes should be obtaining the UB, which in our case is true. AD determines the complexity of an sbox’s ANF, the higher the degree is the more complex it is to launch an algebraic attack on a cipher. Sboxes like those of Ascon/ ISAP having an AD $=2$ are referred to as quadratic and those having AD $=3$ are known as cubic. The functions that posses and AD $=1$ are known as affine.
Side Channel Metrics
During design and analysis phase, DPA SNR and TO of sbox are the two metrics that can be used to determine a cipher’s resistance towards SCA (DPA) attacks. Sboxes that have a high NL and low DU are prone to SCA attacks and have a high resistance towards differential and linear cryptanalysis. Similarly, the sboxes having low values of DPA SNR and TO are resistant towards DPA attacks and vulnerable to differential and linear cryptanalysis. Hence, finding the ideal balance between the two categories of metrics is crucial to the security of a block cipher. Ascon/ ISAP is attaining the almost ideal bounds in terms of NL and DU. Hence, the results of TO and DPA SNR, confirm the notion that Ascon/ ISAP sbox is leaning more on the side of security against linear and differential cryptanalysis. Rest of the sboxes also seem to be attaining UBs on both of SCA metrics.
It is important to note here, that sbox is, although major, yet one mathematical component of a cipher. Designers can always rely on some other design components to balance out the security aspect if using an optimally secure sbox comes at the cost of performance.
Table 5: Results of S-box Analysis
S# Properties Romulus $8\times 8$ Skinny-$128$ Sbox ASCON/ ISAP $5\times 5$ Sbox Elephant $4\times 4$ Sbox GIFT-COFB $4\times 4$ GIFT Sbox Photon-Beetle $4\times 4$ Present Sbox 1 Bijectivity 2 Balancedness 3 Permutation 4 $\mathsf{OP}$ $140$ $26$ $13$ $7$ $9$ 5 $\mathsf{FP}$ $1$ 00 00 00 00 6 $\mathsf{OFP}$ 00 00 $1$ $1$ $1$ 7 $\mathsf{BIC}$ $1.000000$ $1.000000$ $1.000000$ $1.000000$ $1.000000$ 8 $\mathsf{SAC}$ $1.000000$ $1.000000$ $1.000000$ $1.000000$ $1.000000$ 9 $\barwedge_{S}$ $256$ $32$ $16$ $16$ $256$ 10 $\doublebarwedge_{S}$ $4194304$ $8192$ $1024$ $1024$ $1024$ 11 $\xi$ $0.25000$ $0.25000$ $0.25000$ $0.25000$ $0.25000$ 12 $\mathsf{NL}$ $64$ $8$ $4$ $4$ $4$ 13 $\mathsf{BN_{L}}$ $2$ $3$ $2$ $2$ $2$ 14 $\mathsf{LS}$ $601$ $91$ $9$ $9$ $9$ 15 $\mathsf{CI}$ 00 00 00 00 00 16 $\mathsf{DU}$ $64$ $8$ $4$ $4$ $6$ 17 $\mathsf{BN_{D}}$ $2$ $3$ $3$ $2$ $3$ 18 $\mathsf{PC}$ 00 00 00 00 00 19 $\mathsf{UDB}$ $258$ $35$ $3$ $3$ $6$ 20 $\beta\mathsf{U}$ $16$ $16$ $16$ $16$ $256$ 21 $\mathsf{DLU}$ $\frac{64}{128}=0.5$ $\frac{8}{16}=0.5$ $\frac{4}{8}=0.5$ $\frac{4}{8}=0.5$ $\frac{4}{8}=0.5$ 22 $\mathsf{AD}$ $6$ $2$ $3$ $3$ $3$ 23 $\mathsf{DPASNR}$ $6.312455$ $3.015113$ $2.398501$ $2.128608$ $2.398501$ 24 $\mathsf{TO}$ $7.174510$ $4.258065$ $3.266667$ $3.533333$ $3.466667$

### v3

@@ 3.3 Differential Cryptanalysis
Differential cryptanalysis is a technique that relies on exploiting chosen plaintext pairs and resulting ciphertext pairs, in order to gain the information of secret key. The pairs of plaintext and ciphertext are referred as differentials. And these differentials when combined for multiple rounds become differential characteristics of a cipher under study [59]. The probabilities of occurrence of certain output (ciphertext) differences corresponding fixed input (plaintext) differences are computed in the form of DDT. A low value of differential uniformity, high differential branch number, high value of propagation characteristics, and no undisturbed bits are a requirement for an S-box to resist against differential cryptanalysis and its variant attacks.
@@
Where $\varkappa,\varkappa^{\prime}\in F_{2}^{n}$, $\flat,\flat^{\prime}\in F_{2}^{m}$. Ideally, the probability that a particular output difference $(\triangle\flat)$ occurs given a particular input difference $(\triangle\varkappa)$ is $\frac{1}{2^{n-1}}$. Same as LAT, the lookup table of DDT is also of the dimension $2^{n}\times 2^{m}$. Mathematically:-
@@
When an affine transformation is applied to $S$, DDT preserves the maximum entry and weight of coefficients in rows and columns, however their position is subject to a change [117, 118]. Table 4 illustrates bounds on the possible values a single input difference can map to against a single output difference, except for $\triangle_{\varkappa}=0$. Notably, the sum of all entries for a given $\triangle_{\varkappa}$ is upper bounded by $2^{n}$.
Definition 4.
@@
For an S-box to resist differential cryptanalysis a smaller value of DU is required. If an affine transformation is applied to $S$. DU stays invariant [118].
@@
An S-box having DBN greater than $2$ increases the chances of having more active S-boxes in iterative ciphers. DBN does not stay invariant under affine equivalence [112].
@@
$PC(l)$ of $S$ is not variant when an affine transformation is applied [101]. An S-box satisfies PC when $ACT_{max}=0$. For an S-box to satisfy SAC, complying with PC is a pre-requisite [88]. S-boxes having low order $PC(l)$ have weak diffusion characteristics [87].
@@
For a specific nonzero input difference $\triangle_{\varkappa}\in F_{2}^{n}$ of an S-box $S:F_{2}^{n}\rightarrow F_{2}^{m}$, some of the bits in the corresponding output difference $\triangle_{\rho}$ remain unchanged, these specific bits are referred to as undisturbed bits (UDB) [122].
@@
$\triangle_{\flat}=\{\triangle_{\rho}=\rho_{j}\in F_{2}^{m}\ |Pr_{S}(\triangle_{\varkappa}\rightarrow\triangle_{\rho})>0\}$ (24)
Here j=(0,1,2,.m1)j=(0,1,2,....m-1), if $\rho_{j}=d$, when $d\in F_{2}$ is fixed and $\triangle_{\rho}\in\triangle_{\flat}$ then $S$ has UDBs. UDB is invariant under affine transformation. An S-box should have no UDB in order to avoid the construction of truncated differentials. Currently, there is no theoretically proven upper bound on UDB.
@@ 3.4 Boomerang Cryptanalysis
Boomerang attack is a technique of block cipher cryptanalysis based on the similar principle as differential cryptanlysis. However, instead of drawing one differential characteristic for the entire cipher, boomerang attack divides the cipher into two sub ciphers and generates two differentials characteristics also known as boomerang distinguishers. A low value of boomerang uniformity makes an S-box resistant towards boomerang attacks.
@@
The highest value in BCT is known as the boomerang uniformity (BU); where $\triangle_{\varkappa},\nabla_{\flat}]\neq 0 [$67].
@@
BCT and BU of an S-box are invariant under affine transformation [88, 67].
@@ 3.5 Differential-Linear Cryptanalysis
Bar-On et al [69] proposed Differential Linear Connectivity Table (DLCT) in order to study the dependencies between two parts of the cipher so that better differential and linear attack characteristics could be drawn. The DLCT of $S$ is a $2^{n}\times 2^{m}$ table, where rows contain input differences and columns contain output bit masks. For $\triangle_{\varkappa}\in F_{2}^{n}$ and $\rho\in F_{2}^{m}$, the DLCT entry $(\triangle_{\varkappa},\rho)$ can be computed mathematically as follows:-
$\mathsf{DLCT}_{S}(\triangle_{\varkappa},\rho)=\#\{\varkappa\in F_{2}^{n}\mid\ \rho\cdot S(\varkappa)=\rho\cdot S(\varkappa+\triangle_{\varkappa})\}-2^{n-1}$ (27)
@@
[…]
@@
Under affine equivalence, DLCT and DLU are invariant [125, 92]. In order to resist against differential linear attack, an S-box should have low DLU.
@@ 3.6 Algebraic Cryptanalysis
The basic idea behind algebraic attacks involves converting the entire block cipher round function into a system of equations and then trying to solve these equations in order to deduce the secret key. Both, non-linear and linear components are transformed into polynomials. Unlike the linear component, the number of variables for nonlinear layer increases after every next round. Optimal S-boxes are better resistant towards algebraic crypltanalysis if they have a high value of algebraic degree and algebraic immunity, and high degree of interpolation polynomial.
@@
Where the coefficients $c_{i}\in F_{2}$ form the elements of the truth table for ANF of $f(\varkappa)$. An n-bit vectorial boolean function $S:F_{2}^{n}\rightarrow F_{2}^{m}$ has $f_{j}=f_{1},f_{2},f_{3},.....f_{n}$ coordinate functions and each can be uniquely represented in the form of ANF.
@@
The AD is an affine invariant property of any mapping $S$ [115, 88]. S-boxess having a low AD are vulnerable to higher order differential attacks. AD has an inverse relation with CI, the higher is the order of CI, the lower is AD.
@@
In 1996, Jakobsen and Knudsen introduced interpolation attack on block ciphers [126]. In this attack the round function or S-boxes are represented in the form of algebraic expressions. In either case, Lagrange interpolation technique is used to evaluate the algebraic expression. The Lagrange polynomial is constructed using as many plaintext/ciphertext pairs as the number of unknown coefficients in the algebraic representation. Attack complexity depends on number of terms or degree of the resulting polynomial expression.
For any vectorial boolean mapping $S:F_{2}^{n}\rightarrow F_{2}^{m}$ with input $\varkappa_{i}=\{\varkappa_{1},.....,\varkappa_{n}\}$ and output $\flat_{i}=\{\flat_{1},.....,\flat_{n}\}$ where i=1,,ni=1,......,n, we find the Interpolation Polynomial (IP) $P_{i}(\varkappa)$ as follows:
@@
IP is invariant under affine equivalence [127], and a low degree of IP makes any S-box susceptible to interpolation attack.
@@ 3.7 Side Channel Analysis
Attacks on cryptographic implementations have evolved significantly since Kocher’s first attempt to use physical leakages from a cryptosystem to extract sensitive information. Due to the increasing threat, designers are always trying to come up with efficient ways to resist against side channel attacks [128]. Starting from the design of components such as S-boxes with good cryptographic properties to resist side channels and moving up to the overall cipher level by employing countermeasures such as boolean or polynomial masking. Properties such as DPA SNR and transparency order are nowadays actively used to determine an S-box’s resilience against DPA attacks. However, striking a perfect balance between the metrics related to traditional cryptanalysis and side channel analysis is an open research area. Currently, designers try to achieve cryptographic profiles that are well suited with the requirements of a particular infrastructure or standardization process (as in NIST LWC). In ideal terms, for an S-box to be resistant towards attacks such as DPA, it needs to posses a low DPA SNR and transparency order.
@@
If an S-box has small value of TO, it is considered highly resistant against DPA attacks [73]. When affine transformation is applied to an S-box, TO does not stay invariant [95].
4 Discussion and Results
We have given the details of our findings on Romulus, ASCON/ ISAP, Elephant, Photon-Beetle, and GIFT-COFB in Tables 6-7 against dedicated cryptographic profiles. We briefly explain them in subsequent subsections.
4.1 Diffusion Metrics
@@
If we look at the basic avalanche characteristics of all these S-boxes they satisfy the criteria of balancedness. This helps reduce the chances of getting any high probability linear approximations, in turn assuring the resistance of these S-boxes towards linear cryptanalysis [93]. The S-boxes under analysis are bijective, so they satisfy the criteria of permutation. From the bounds discussed in Table 4, we have developed an understanding that ideally the S-boxes should have a full cycle length for order of permutation. Presumably, S-boxes having short iterative periods are vulnerable to attacks. In order to avoid such attacks designers can use affine equivalence to improve the order of permutation of these S-boxes [62, 63, 97, 98].
Out of the six, Ascon/ ISAP S-box has a FP at $\varkappa=255$. Photon-Beetle/ Elephant/ GIFT-COFB, each has one OFPs at $\varkappa=14/8/3$. S-boxes having FPs can alter the overall diffusion power that an iterative cipher has to offer and make it vulnerable to invariant subspace attacks, it may be rectified using affine equivalent S-box.
The upper bounded BIC value for all the S-boxes indicates that the input and output avalanche vectors are fully correlated.
Subsequently, none of the S-box satisfies the principles of completeness and avalanche depicted by their SAC failure. The AI values for all the S-boxes achieve the upper bound, and the values for that of SSI seem a bit better but do not satisfy the lower bound requirement. As discussed earlier, for an S-box to satisfy SAC, one of the prerequisites is compliance with PC. In Tables 6-7 we can see that all S-boxes have PC =0=0, which again points towards the weak avalanche profile. These findings indicate that an attacker can utilize this information to mount statistical attacks on the underlying cipher [99, 129].
4.2 Linear Metrics
@@
Security towards linear cryptanalysis attacks is ensured when an S-box has a low LAP and higher non linearity. Our results show that all $6$ S-boxes under analysis have achieved the desired bounds for both properties. The LBN of all $6$ S-boxes is closer to the LB, which is opposite to the ideal requirements. However, if the design requires it, LBN can be improved using affine transformations. The presence of LS makes it possible for the cipher to leak significant amount of information if specific group(s) of bits in the input are flipped. All the S-boxes are first order correlation immune, which makes them secure against zero-correlation attacks.
4.3 Differential and Boomerang Metrics
@@
Almost Perfect Nonlinear (APN) S-boxes, in odd dimensions are differentially $2$ uniform [63]. The only odd dimension case in S-boxes under study is of Ascon/ ISAP, and since its has a DU: $8$. The remaining S-boxes have a DU within the desired bounds. The DBN of Elephant and Photon-Beetle S-box satisfies the UB, Ascon/ ISAP is closer to the UB, and for Romulus and GIFT-COFB DBN value stands on the LB. A good DBN ensures more number of active S-boxes across multiple rounds in block ciphers. The UDB are detected in all $6$ S-boxes. In [122] authors have extensively described how the presence of UDB in an S-box impacts other cryptographic properties, such as the occurrence of LS, as we can see in Tables 6-7.
Tezcan et al. used these UDB and LS to devise 4 and 5 round impossible, truncated and improbab differentials, and break 5 out of 12 Ascon rounds [130].
In regards to boomerang cryptanalysis, the BU for all S-boxes is attaining the UB, which is opposite to the required bounds in order to be secure against Boomerang attacks.
4.4 Differential-Linear Metrics
@@
Combined attacks make it easier for an attacker to break a cipher by defining short differential and linear characteristics, as in the case of differential-linear cryptanalysis. Hence, attaining the LB on DLU is crucial to gain resistance against these attacks. In case of S-boxes under analysis, we can observe from Tables 6-7, that the DLU is a little way up from the LB, hence paving a way for differential-linear attack with a reduced complexity.
4.5 Algebraic Metrics
@@
Previously, we have talked about the inverse relation between CI and AD, and we observed that all $6$ S-boxes have a CI =0=0. So given the inverse relation, AD of all S-boxes should be obtaining the UB, which in our case is true. AD determines the complexity of an S-box’s ANF, the higher the degree is the more complex it is to launch an algebraic attack on a cipher. S-boxes like those of Ascon/ ISAP having an AD =2=2 are referred to as quadratic and those having AD =3=3 are known as cubic. The functions that posses and AD =1=1 are known as affine.
4.6 Side Channel Metrics
During design and analysis phase, DPA SNR and TO of S-box are the two metrics that can be used to determine a cipher’s resistance towards SCA (DPA) attacks. From the conventional cryptanalysis perspective, S-boxes having a high NL and low DU are prone to SCA attacks and have a high resistance towards differential and linear cryptanalysis, and vice versa. Hence, finding the ideal balance between the two axes of attack metrics is crucial to the security of a lightweight block cipher. Since the S-boxes of LWC finalists vary in size (4, 5 and 8 bits), a direct comparison of the results we obtained of their side-channel metrics is not fair, since the required optimal bounds for each size are different (see Table 4). Hence to draw a meaningful comparison, we normalize the DPA SNR and TO using the ideal or worse case bounds as a reference point.
  • DPA SNR can be normalized using the identity-case bound [72], as follows:
    $\mathsf{DPA_{SNR}(I)=\sqrt{m}}$ (36)
    Here, in the case of a balanced S-box $m$ is the number of input/output bits. We now define the normalized DPA SNR as follows:-
    $\mathsf{DPA_{SNR}(norm)=\frac{{DPA_{SNR}(S)}-{DPA_{SNR}(I)}}{DPA_{SNR}(S)}}$ (37)
    The results of the suggested normalization are in the range $[0,1]$. Here, a value 0 indicates that the S-box under analysis has a DPA SNR equal to that of an identity S-box, which theoretically is the best DPA SNR possible for an S-box to be resistant towards DPA attacks. For example, in the case of 8-bit S-box, the identity has a DPA SNR $2.82$ and the DPA SNR of Romulus is $6.312455$. The normalized DPA SNR in the 8-bit case is $0.552$, suggesting that Romulus S-box potentially leaks $55.2\%$ more than the identity S-box. In the case of 4-bit sbox, the identity has a DPA SNR $2$, while GIFT, PHOTON and Elephant have DPA SNR $2.128608$ and $2.398501$, respectively. The normalized DPA SNR in the case of GIFT is $0.064304$, and in the case of PHOTON and Elephant is $0.1992505$, respectively. The normalization in the case of 4-bit S-boxes suggests GIFT has the lowest leakage amongst the three 4-bit S-boxes, as it leaks only $6.4304\%$ more than the 4-bit identity S-box.
  • TO can be normalized using the worse-case (highest leakage) bound [73], as follows:-
    $\mathsf{{TO}_{highest-leakage}=m}$ (38)
    Here, in the case of a balanced S-box $m$ is the number of input/output bits. Based on the worse-case bound, we now define the normalized TO as follows:-
    $\mathsf{TO(norm)=\frac{TO_{highest-leakage}-{TO_{S}}}{TO_{highest-leakage}}}$ (39)
    The results of this normalization lie in the range [0,1], where a factor 0 depicts that the TO is the same as the worst-case scenario, i.e., the maximum TO value, implying the highest leakage in the case of DPA attacks. Any value greater than 0 suggests relative improvement in the TO value and hence reduced leakage. For example, in the case of 8-bit S-box, the worse-case TO is $8$ and the TO of Romulus is $7.174510$. The normalized TO in the 8-bit case is $0.1031$, suggesting that Romulus S-box potentially leaks $10.31\%$ less than the worse-case S-box.
5 Exploring Trade-offs in S-box Design
Based on our analysis in Section 4, several finalist S-boxes exhibit weaknesses in classical cryptographic metrics — such as fixed points, differential uniformity, non-linearity, linear structures, etc. So to explore the possibility of improving the cryptographic robustness of S-boxes beyond those used in the NIST LWC finalists, we generated a new $4\times 4$ bit test S-box using the PEIGEN [75] tool for S-box evaluation and generation, as shown in Table 8, using the following constraints:-
  • Permutation: ✓
  • Differential Uniformity: 4
  • Algebraic Degree: 3
  • Linear Structures: 0
  • Implementation Cost: 8
In PEIGEN, cost is a composite metric that captures the implementation efficiency in terms of both hardware and software, focusing on the following traits:-
  • Gate Equivalent (GE): refers to a 2-input NAND gate unit,
  • Bitslice Gate: for software implementations (optimized 32-bits),
  • Circuit depth in terms of latency (depth vs area), and
  • Multiplicative complexity: minimizing number of non-linear operations (for side channel secure ciphers where masking and other countermeasures are involved).
Table 8: Proposed 4-bit S-box
4-bit 0 4 5 1 c 9 6 7 3 a e 8 b 2 f d
Fixing the implementation cost ensures that the resultant S-box implementation lies within the acceptable practical bounds for lightweight cryptography implementations, while retaining the desired cryptographic profile.
5.1 Comparison and Analysis
Table 7, highlights the comparison between the S-box new 4-bit S-box and and the 4-bit S-boxes from GIFT-COFB, Photon-Beetle and Elephant. From the classical cryptanalysis perspective all four S-boxes have the same differential uniformity, non-linearity, linear approximation, algebraic degree and correlation immunity. A noticeable difference occurs in the properties of the new S-box when compared to the other three, as follows:-
- The values for BIC and SAC are better.
- No linear structures and undisturbed bits.
- Boomerang uniformity has improved.
Some of the trade-offs that exist in the properties of the new S-box are as follows:-
- $3$ fixed points are present in the new S-box while the others have null.
- DPA-SNR value of $2.5786$ is a little higher than for all three finalists, indicating slightly weaker resistance to side-channel (DPA) leakage, but offers optimal resistance against conventional cryptanalysis.
- TO of $3.4$ is a little higher than that of Elephant ($3.26$), but a little lower than that of PHOTON and GIFT-COFB, indicating a reasonable balance between leakage and classical security.
The current findings underscore that even when the classical cryptanalysis properties are optimized within practical bounds, striking a perfect balance with the side-channel metrics may not be possible. In practice, identifying such balanced S-boxes remains a trial-and-error process that involves testing multiple generation strategies and parameter choices until an acceptable trade-off is found. Moreover, it is important to emphasize that the S-box alone does not define the overall security of a lightweight block cipher. Resistance to differential and linear cryptanalysis, as well as to side-channel analysis, also depends heavily on other design aspects such as the diffusion layer, round constants, and key schedule, etc. A strong S-box is therefore only one building block of a well-rounded cryptographic primitive.
6 Scope and Limitations
Our research presents a comprehensive evaluation of the cryptographic properties of S-boxes used in finalist ciphers of the NIST LWC competition. For the sake of clarity and accuracy it is important that we acknowledge the scope and limitations of our analysis.
Firstly, we would like to clarify that our analysis of S-boxes is conducted from a theoretical and metric-based perspective. We do not claim to perform any practical side-channel attacks (e.g., DPA, SPA) or cryptanalytic attacks (e.g., differential, linear) on the full/reduced-rounds structure of NIST LWC finalists.
Strong S-boxes contribute significantly to overall cipher security. We have evaluated the S-boxes individually, based on well-established cryptographic metrics listed in Table 3. Such compact analysis is a crucial first-step in determining the strength of the confusion layer when designing a new cipher, but it should be noted that full cipher-level cryptanalysis and side-channel analysis requires holistic evaluation – involving linear layers, modes of encryption, masking strategies, implementation environments, and leakage models. These are beyond the scope of this study and can be seen as potential future work.
Therefore, we specifically refrain from extrapolating our S-box component-level findings to make any assertions on the overall security strength of the cipher.